OpenStudy (anonymous):

Given that f(x)=x7+5x+1 and g(x)=7x+8, find [fogof](x) and (f+3g)(-1)

OpenStudy (anonymous):

well, why not start calculating

zepdrix (zepdrix):

I prefer the other notation, it's a little easier to work with.\[\Large\bf\sf \left[f\circ g\circ f\right](x)\quad=\quad f\left[g(f(x))\right]\]So let's work from the outside first I guess,\[\Large\bf\sf f\left[\color{royalblue}{x}\right]\quad=\quad (\color{royalblue}{x})^7+5(\color{royalblue}{x})+1\]

zepdrix (zepdrix):

\[\Large\bf\sf f\left[\color{royalblue}{g(f(x))}\right]\quad=\quad (\color{royalblue}{g(f(x))})^7+5(\color{royalblue}{g(f(x))})+1\]

zepdrix (zepdrix):

Hmm ya this is gonna get ugly :O

zepdrix (zepdrix):

So I'm plugging f into g here, \[\Large\bf\sf g(\color{orangered}{x})=7(\color{orangered}{x})+8\qquad\implies\qquad g(\color{orangered}{f(x)})\quad=\quad 7\color{orangered}{f(x)}+8\]Then let's plug f(x) in for the orange part,\[\Large\bf\sf g(\color{orangered}{f(x)})\quad=\quad 7(\color{orangered}{x^7+5x+1})+8\]And now we plug this whole thing in for the blue stuff from earlier.

zepdrix (zepdrix):

\[\bf\sf f\left[\color{royalblue}{g(f(x))}\right]\quad=\quad (\color{royalblue}{7(x^7+5x+1)+8})^7+5(\color{royalblue}{7(x^7+5x+1)+8})+1\]

zepdrix (zepdrix):

I don't know if the colors are actually helping on this problem.. there are so many layers :( Kinda gets muddled up and hard to read after a while.

zepdrix (zepdrix):

But umm.. you certainly would not want to simplify that... that's for sure :\ Just stop there for the first one... :p

zepdrix (zepdrix):

I dunno, you could simplify it a little bit. But definitely don't expand out the 7th power. That would be crazy +_+

OpenStudy (anonymous):

that's the method zepdrix

zepdrix (zepdrix):

What's what method?

OpenStudy (anonymous):

the simplification of the above question. thanks

OpenStudy (anonymous):

He's got it. Catch you later with a more complicated math equation...

zepdrix (zepdrix):

Oh you were just testing us or something? XD lol